# AI Agentic Vulnerability Injection And Transformation with Optimized Reasoning

### v1

@@ Abstract
The increasing complexity of software systems and the sophistication of cyber-attacks have underscored the critical need for effective automated vulnerability detection and repair systems. Traditional methods, such as static program analysis, face significant challenges related to scalability, adaptability, and high false-positive and false-negative rates. AI-driven approaches, particularly those using machine learning and deep learning models, show promise but are heavily reliant on the quality and quantity of training data. This paper introduces a novel framework designed to automatically introduce realistic, category-specific vulnerabilities into secure C/C++ codebases to generate datasets. The proposed approach coordinates multiple AI agents that simulate expert reasoning, along with function agents and traditional code analysis tools. It leverages Retrieval-Augmented Generation for contextual grounding and employs Low-Rank approximation of weights for efficient model fine-tuning. Our experimental study on 116 code samples from three different benchmarks suggests that our approach outperforms other techniques with regard to dataset accuracy, achieving between 89% and 95% success rates in injecting vulnerabilities at function level.
Index Terms:
Automated Vulnerability Detection, AI-Driven Security, Agentic Workflows
@@ I Introduction
Software vulnerabilities refer to weaknesses in system security requirements, design, implementation, or operation, that could be accidentally triggered or intentionally exploited, resulting in a violation of the system’s security policy [1]. The rapid growth in software complexity, coupled with the sophistication of cyber-attacks, poses a significant threat to the global security and stability of digital infrastructures. In 2024 alone, the total number of publicly reported vulnerabilities increased by 25% [2].
Consequently, effective automated vulnerability detection and repair systems have become critically important. Traditional methods, such as static program analysis, face significant challenges related to scalability, adaptability, and high false-positive and false-negative rates [3] [4]. Due to their reliance on manually crafted rule-based heuristics and the combinatorial complexity of their analysis, these programmatic approaches inherently constrain opportunities for potential significant breakthroughs.
AI-driven approaches for code vulnerability detection [5, 6, 7, 8, 9, 10, 11] and repair [12, 13, 14] have recently attracted considerable attention in the research community. Data-driven methods, especially those using machine learning and deep learning models, show promise in identifying and addressing a wide range of vulnerabilities. However, like most AI-driven systems, their effectiveness heavily relies on the quality and the quantity of the training data [11]. This dependency has become even more critical with the emergence of Large Language Models (LLMs), which can have billions of parameters to train. Nevertheless, the scarcity of large, high-quality training datasets has become a major obstacle to the progress of these techniques [11] [6] [5]. Existing vulnerability datasets can be grouped into the following categories based on how the code samples were collected and annotated:
@@
  • Large realistic datasets: Several widely used large-scale vulnerability datasets, including ReVeal [5], BigVul [17], CVEfixes [18], CrossVul [19] and Diversevul [10] rely on heuristic-based labeling functions. These approaches typically involve identifying vulnerability-fixing commits from public security databases, such as the National Vulnerability Database (NVD), and labeling the pre-commit versions of affected functions as vulnerable. While these datasets contain code samples from real-world production software, the reliance on heuristics introduces noise and label inaccuracies. Studies have reported label accuracies as low as 25% for some datasets [11] [20]. Additionally, datasets like D2A [21] and Draper [22] use static analyzers to label code, which results in numerous false positives due to the limitations of static analysis tools. The presence of noisy or incorrect labels in all of these datasets poses challenges for training reliable vulnerability detection models [10] [11].
@@
  • Reliable labels datasets: To enhance label accuracy, some datasets employ manual expert annotation or stricter heuristics. For instance, SVEN [23] manually labels 1,606 functions, improving label quality but limiting scalability due to resource constraints and coverage of only 9 types of Common Weakness Enumeration (CWE). Alternatively, PrimeVul [11] applied rigorous heuristics, claiming human-level labeling accuracy across  7,000 vulnerable and  229,000 benign functions spanning over 140 CWEs. Despite these advancements, the creation of larger high-quality datasets remains challenging to scale, especially when corresponding pairs of vulnerable and benign code samples are required.
@@
  • Vulnerability injection-based datasets: Several approaches for injecting bugs into production software have been proposed and implemented [24, 25, 26, 27]. Methods such as VULGEN [28] and VinJ [29] combine pattern mining with deep learning-based localization to identify injection sites and apply vulnerability-inducing edits. These approaches scale well and produce diverse, realistic samples. However, injected bugs may lack the semantic and contextual richness of real-world vulnerabilities, and current techniques typically support only a limited set of CWE types. Moreover, the accuracy of injected datasets remains a concern: for example, VULGEN achieves only a 69% success rate in generating truly vulnerable samples, indicating potential noise and limited reliability.
@@
To address these challenges, we introduce AVIATOR (AI Agentic Vulnerability Injection And Transformation with Optimized Reasoning), an AI agentic workflow designed to automatically introduce realistic, category-specific vulnerabilities into secure C/C++ codebases. AVIATOR coordinates multiple AI agents that simulate expert reasoning, along with function agents and traditional code analysis tools. It leverages Retrieval-Augmented Generation (RAG) [30] for contextual grounding and employs Low-Rank Adaptation (LoRA) [31] for efficient model fine-tuning.
@@
To evaluate the performance of AVIATOR in introducing vulnerabilities we conducted an experimental study on 116 code samples from three different benchmarks used in the literature [16, 32, 11]. We leveraged formal verification tool ESBMC [33] for supported vulnerabilities and manual verification by experienced computer security researchers. We compare AVIATOR with existing automated large-scale realistic software vulnerability dataset creation technique. The results suggest a potential for the AVIATOR to outperform other techniques with regard to accuracy of dataset labels. It achieved between 89% and 95% success rates in injecting vulnerabilities across the different benchmarks.
@@
  • Agentic Injection Workflow: To the best of our knowledge, we are the first to develop a modular, multi-agent pipeline that replicates the reasoning processes of cybersecurity experts to generate plausible vulnerability injections.
  • Contextual Augmentation: By integrating Retrieval-Augmented Generation (RAG) modules, we ensure that vulnerability transformations are grounded in diverse, real-world code contexts.
  • Efficient Fine-Tuning: We utilize Low-Rank Adaptation (LoRA) to fine-tune large language models on small-scale datasets with minimal computational cost, thereby enhancing scalability.
  • Empirical Validation: Our preliminary evaluation demonstrates that AVIATOR achieves between 89% and 95% success in generating plausible vulnerabilities, facilitating the creation of large-scale, accurately labeled, and diverse datasets for training advanced detection models. Through our experiments, we have established the defining characteristics of high-quality software vulnerability datasets, essential for both training AI models and benchmarking purposes.
The remainder of this paper is organized as follows: Section 2 reviews existing methods for vulnerability detection and data generation. Section 3 details the architecture and agentic workflow of AVIATOR. Section 4 outlines our experimental setup and evaluation metrics. Section 5 presents empirical results and analysis. Section 6 discusses limitations and future work, and Section 7 concludes.
@@ II-A Properties of High-Quality Vulnerability Datasets
  • Diverse: The dataset should cover varied coding styles and vulnerability types to enable generalizable learning and comprehensive benchmarking across scenarios.
@@
  • Paired samples, benign and vulnerable versions, are ideal for repair tasks and contrastive learning.
@@ II-B Motivation
Vulnerability injection presents a promising alternative. By design, it enables scalability through automation and preserves credibility when applied to real-world codebases. However, the effectiveness of this approach depends heavily on the injection technique’s ability to introduce diverse and accurate vulnerabilities. Current state-of-the-art methods remain limited in this regard: recent studies report coverage of only 19 distinct CWE classes and achieve an injection success rate of just 69% [28, 29]. Such limitations constrain both the diversity and reliability of resulting datasets, particularly for benchmarking purposes. Nonetheless, these initial results suggest that with more refined techniques, vulnerability injection has the potential to systematically produce accurate, realistic, diverse, and large-scale datasets that better support the development and test of modern AI-based vulnerability detection and repair systems, as well as traditional static analyzers.
@@ II-C Vulnerability Injection
III Proposed Framework
Existing vulnerability injection techniques continue to face fundamental limitations in both coverage and reliability. These methods generally support only a limited range of vulnerability types and exhibit low success rates in injection. This is largely due to their inability to adapt to the structural and contextual intricacies of real-world code found in production, as well as a lack of deep semantic reasoning.
To overcome these limitations, the proposed AVIATOR framework comprises two core modules: (1) a vulnerability injection module that utilizes retrieval-augmented generation (RAG) and supervised fine-tuning with Low-Rank Adaptation (LoRA) to guide code transformations; and (2) a validation module that employs LLM-based discriminators and a static analyzer to meticulously verify the presence and accuracy of the injected vulnerabilities. These two modules are illustrated in Figure 1.
In the following, we present the AVIATOR framework’s core components and methodologies. We first define AI agentic workflows, where specialized agents collaborate through structured execution paths to enhance modularity and error correction, mimicking cybersecurity experts’ analytical processes. We then detail the Vulnerability Injection Module of the workflow, which transforms secure code into vulnerable code using semantic analysis and advanced techniques like retrieval-augmented generation and fine-tuning. Additionally, we discuss the Validation Module, which ensures the accuracy and presence of injected vulnerabilities through a combination of deterministic tools and language model agents.
@@ III-A AI Agentic Workflows
Agentic workflows are systems in which a collection of agents are orchestrated to perform a complex task by following a predefined execution path [35]. Each agent is responsible for a specific action and transforms structured inputs into outputs that serve subsequent agents. Agents can be instantiated as large language models (LLMs) or deterministic programmatic tools (e.g., static analyzers).
@@
Formally, we define an agentic workflow $W_{k}$ of $k$ agents as a directed execution graph, where each node represents an agent and edges denote conditional transitions. Each agent $A_{i}$ transforms structured inputs into outputs, given by $A_{i}:X_{i}\times E\mapsto Y_{i}$, with $E$ denoting external contextual artifacts (e.g., retrieved knowledge). Transitions depend on conditions evaluated dynamically from prior outputs.
@@
This architectural paradigm offers advantages over monolithic reasoning or linear pipelines. First, it enables modular decomposition: the overall task is split into smaller, logically coherent subtasks aligned with the competencies of individual agents. This reduces the reasoning burden on any single component and improves interpretability. Second, it allows structured context propagation: each agent receives only the relevant inputs for its subtask, while outputs from earlier agents can be enriched with retrieval or validation results. Third, agentic workflows support corrective execution: if an agent detects inconsistencies or failures in prior outputs, the workflow can re-invoke earlier agents to revise decisions or recover from errors.
In AVIATOR, we implement an agentic workflow that mirrors the step-by-step reasoning of a cybersecurity analyst. The system orchestrates specialized agents responsible for tasks such as vulnerability selection, injection point identification, transformation synthesis, and post-injection validation, as detailed in III-B, III-C. This modular design promotes targeted decision-making, minimizes error propagation, and ensures that vulnerability injection proceeds in a systematic, verifiable, and context-aware manner.
III-B Vulnerability Injection Module
The Vulnerability Injection Module in AVIATOR plays a crucial role by converting a benign code snippet into a version that includes a specific vulnerability. Central to the framework, this module ensures that the injected vulnerability is not only present but also realistic, minimal, and contextually coherent. It comprises the following components.
III-B1 Semantic Analysis and Injection Point Identification

### v2

@@ Abstract
The increasing complexity of software systems and the sophistication of cyber-attacks have underscored the critical need for effective automated vulnerability detection and repair systems. Data-driven approaches using deep learning models show promise but critically depend on the availability of large, accurately labeled datasets. Yet existing datasets either suffer from noisy labels, limited range of vulnerabilities, or fail to reflect vulnerabilities as they occur in real-world software. This also limits large-scale benchmarking of such solutions. Automated vulnerability injection provides a way to directly address these dataset limitations, but existing techniques remain limited in coverage, contextual fidelity, or injection success rates. In this paper, we present AVIATOR, the first AI-agentic vulnerability injection workflow. It automatically injects realistic, category-specific vulnerabilities for high-fidelity, diverse, large-scale vulnerability dataset generation. Unlike prior monolithic approaches, AVIATOR orchestrates specialized AI agents, function agents and traditional code analysis tools, that replicate expert reasoning. It combines semantic analysis, injection synthesis enhanced with LoRA-based fine-tuning and Retrieval-Augmented Generation, as well as post-injection validation via static analysis and LLM-based discriminators. This modular decomposition allows specialized agents to focus on distinct tasks, improving robustness of injection and reducing error propagation across the workflow. Evaluations across three distinct benchmarks demonstrate that AVIATOR achieves 91%–95% injection success rates, significantly surpassing existing automated dataset generation techniques in both accuracy and scope of software vulnerabilities.
@@ I Introduction
The rapid growth in software complexity, coupled with the sophistication of cyber-attacks, poses a significant threat to the global security and stability of digital infrastructures. In 2024 alone, the total number of publicly reported vulnerabilities rose by 25% [1]. Software vulnerabilities refer to weaknesses in system security requirements, design, implementation, or operation, that could be accidentally triggered or intentionally exploited, resulting in a violation of the system’s security policy [2]. These trends highlight the urgent need for scalable and effective automated vulnerability detection and repair systems. Traditional methods, such as static program analysis, have played an important role. However, their reliance on manually crafted heuristics and the combinatorial complexity of their analysis makes them fundamentally limited, with persistent challenges in scalability, adaptability, and false alarm rates [3, 4].
In response, AI-based approaches for code vulnerability detection [5, 6, 7, 8, 9, 10, 11] and repair [12, 13, 14] have recently attracted considerable attention in the research community. These data-driven techniques, show strong potential in identifying and addressing a wide range of vulnerabilities. However, their effectiveness heavily relies on the quality and scale of the training data [11, 15]. This dependency has become even more critical with the emergence of Large Language Models (LLMs) with billions of parameters to train. The scarcity of high-quality, large-scale datasets therefore remains a central obstacle to advancing automated software security [11, 6, 5]. Existing vulnerability datasets can be grouped into the following categories based on how the code samples were collected and annotated:
@@
  • Large realistic datasets: Several widely used large-scale vulnerability datasets, including ReVeal [5], BigVul [18], CVEfixes [19], CrossVul [20] and Diversevul [10] rely on heuristic-based labeling functions. These approaches typically involve identifying vulnerability-fixing commits from public security databases, such as the National Vulnerability Database (NVD) [21], and labeling the pre-commit versions of affected functions as vulnerable. While these datasets contain code samples from real-world production software, the reliance on heuristics introduces noise and label inaccuracies. Studies have reported label accuracies as low as 25% for some datasets [11] [22]. Additionally, datasets like D2A [23] and Draper [24] use static analyzers to label code, which results in numerous false positives due to the limitations of static analysis tools. The presence of noisy or incorrect labels in all of these datasets poses challenges for training reliable vulnerability detection models [10] [11].
@@
  • Reliable labels datasets: Some datasets attempt to improve label accuracy through manual expert annotation or by enforcing very strict heuristics. For instance, SVEN [25] manually labels 1,606 functions, which ensures clean labels, but covering only 9 Common Weakness Enumeration (CWE) types. PrimeVul [11] applies rigorous heuristics, claiming human-level labeling accuracy across  7,000 vulnerable and  229,000 benign functions spanning over 140 CWEs. Yet, manual curation is inherently tedious and produces only small datasets, while overly strict heuristics hinder scalability.
@@
  • Vulnerability injection-based datasets: Several approaches for injecting bugs into production software have been proposed and implemented [26, 27, 28, 29]. Methods such as VULGEN [30], VinJ [31] and VGX [15] combine pattern mining with deep learning-based localization to identify injection sites and apply vulnerability-inducing edits. These approaches scale better and produce realistic samples, but they remain constrained in both coverage and reliability. In practice these, injected bugs are limited to single-statement edits and may lack the semantic and contextual richness of real-world vulnerabilities. Moreover, the accuracy of injected datasets remains a concern: for example, VULGEN achieves only a 69% success rate in generating truly vulnerable samples, indicating potential noise and limited reliability.
@@
To address these challenges, we introduce a novel AI agentic workflow designed to automatically inject realistic, diverse, category-specific vulnerabilities for large-scale generation of high-quality vulnerability datasets. This framework, called AVIATOR (AI Agentic Vulnerability Injection And Transformation with Optimized Reasoning) coordinates multiple AI agents - simulating expert reasoning - alongside functional agents and traditional code analysis tools. It was designed by understanding the full process of manual injection and splitting it into elementary subtasks, relying on tools and self-correction loops to provide optimal guidance to LLMs. Its design was validated by three cybersecurity experts (including two co-authors). To further boost performance, it leverages Retrieval-Augmented Generation (RAG) [32] for contextual grounding and employs Low-Rank Adaptation (LoRA) [33] for efficient model fine-tuning. Crucially, AVIATOR requires no pattern mining, freeing it from the rigid constraints of mined edits and allowing scalable generation of vulnerabilities that span more CWE categories, with greater semantic and contextual realism.
@@
To evaluate the performance of AVIATOR in introducing vulnerabilities we conducted an experimental study on code samples from three different benchmarks used in the literature [17, 34, 11]. We leveraged formal verification tool ESBMC [35] for supported vulnerabilities in smaller functions and manual verification by experienced computer security researchers for the rest. We compare AVIATOR with existing automated large-scale realistic software vulnerability dataset creation technique. The results suggest a potential for the AVIATOR to outperform other techniques with regard to accuracy of dataset labels. It achieved between 91% and 95% success rates in injecting vulnerabilities across the different benchmarks.
AVIATOR is a generic framework capable of supporting injection across a large CWE spectrum. Our current instantiation already covers 133 CWE types, far exceeding VGX (23 CWEs) and demonstrating the diversity potential of datasets generated with AVIATOR. Unlike prior injection-based approaches, AVIATOR does not rely on language-specific parsers or handcrafted pattern mining. Its agentic design instead leverages LLM-based reasoning making it both broadly applicable and adaptable across programming languages.
@@
  • Agentic Injection Workflow: To the best of our knowledge, we are the first to explore vulnerability injection using an AI-agentic workflow. We carefully designed a multi-agent pipeline that replicates the reasoning processes of cybersecurity experts to generate plausible vulnerability injections for large-scale quality dataset.
  • Contextual Augmentation: By integrating Retrieval-Augmented Generation modules, we ensure that vulnerability transformations are grounded in diverse, real-world code contexts.
  • Efficient Fine-Tuning: We utilize Low-Rank Adaptation (LoRA) to fine-tune LLMs on small-scale datasets with minimal computational cost, thereby enhancing scalability.
  • Empirical Validation: Our preliminary evaluation demonstrates that AVIATOR achieves between 89% and 95% success in generating plausible vulnerabilities, facilitating the creation of large-scale, accurately labeled, and diverse datasets Through our experiments, we have established the defining characteristics of high-quality software vulnerability datasets, essential for both training AI models and benchmarking purposes.
The remainder of this paper is organized as follows: Section 2 reviews existing methods for vulnerability detection and data generation. Section 3 details the architecture and agentic workflow of AVIATOR. Section 4 outlines our experimental and evaluation setup. Section 5 reports empirical results and analysis. Section 6 discusses limitations and future work, and Section 7 concludes.
@@ II-A Properties of High-Quality Vulnerability Datasets
  • Diverse: The dataset should cover varied categories of vulnerability and coding styles to enable generalizable learning and comprehensive benchmarking across scenarios.
@@
  • Paired benign and vulnerable samples, are ideal for repair tasks and contrastive learning.
@@ II-B Motivation
Vulnerability injection presents a promising alternative. By design, this data augmentation technique enables scalability through automation and preserves credibility when applied to real-world codebases. However, the effectiveness of this approach depends heavily on the injection technique’s ability to introduce diverse and accurate vulnerabilities. Current state-of-the-art methods remain limited in this regard: recent studies report coverage of only 19 distinct CWE classes and achieve an injection success rate of just 69% [30, 31]. Such limitations constrain both the diversity and reliability of resulting datasets, particularly for benchmarking purposes. Nonetheless, these initial results suggest that with more refined techniques, vulnerability injection has the potential to systematically produce accurate, realistic, diverse, and large-scale datasets that better support the development and benchmarking of modern AI-based vulnerability detection and repair systems, as well as traditional static analyzers.
@@ II-C Vulnerability Injection
III Proposed Framework
Existing vulnerability injection techniques continue to face fundamental limitations in both coverage and reliability. These methods generally support only a limited range of vulnerability types and exhibit low success rates in injection. This is largely due to their inability to adapt to the structural and contextual intricacies of real-world code found in production, as well as a lack of deep semantic reasoning. To overcome these limitations, the proposed AVIATOR framework in figure 1, organizes the injection process as an AI-agentic workflow, explicitly mirroring the reasoning steps that a human security analyst would take.
This section details the design principles underlying AVIATOR, the agentic workflow abstraction on which it is built, and the two core modules of the framework: vulnerability injection and validation.
@@ III-A AI Agentic Workflows
AVIATOR builds on the concept of an agentic workflow, where a complex task is decomposed into smaller, well-defined subtasks executed by specialized agents. Each agent is instantiated as either a large language model (LLM) or a programmatic tool such as a static analyzer. Agents are connected through a directed execution graph with typed inputs and outputs, and conditional transitions that allow for retries or fallbacks when earlier steps fail. In contrast to monolithic program transformations that rely on a single model to perform injection end-to-end, this design constrains the scope of each agent and enables structured feedback between stages.
@@
Formally, an agentic workflow $W_{k}$ with $k$ agents is a directed graph where nodes correspond to agents and edges represent conditional transitions. Each agent $A_{i}$ is defined as a transformation $A_{i}:X_{i}\times C_{i}\mapsto Y_{i}$, where $X_{i}$ is the structured input, $C_{i}$ is an optional set of contextual artifacts (e.g., CWE descriptions, retrieved examples, static analysis reports), and $Y_{i}$ is the output. Transitions are governed by conditions dynamically evaluated from prior outputs, allowing the workflow to re-invoke or bypass agents as needed.
@@
The agentic workflow design provides advantages over monolithic reasoning or linear pipelines. First, it enables modular decomposition: the overall task is split into smaller, logically coherent subtasks aligned with the competencies of individual agents. This reduces the reasoning burden on any single component and improves interpretability. Second, it enables structured context propagation: each agent receives only the relevant inputs for its subtask, while outputs from earlier agents can be enriched with retrieved or validation results. Third, agentic workflows support corrective execution: if an agent detects inconsistencies or failures in prior outputs, the workflow can re-invoke earlier agents to revise decisions or recover from errors. This modular design promotes targeted decision-making, minimizes error propagation, and ensures that vulnerability injection proceeds in a systematic, verifiable, and context-aware manner. These properties make agentic workflows well suited to complex code transformation tasks such as vulnerability injection.
III-B Design Methodology and Iterative Refinement
The design of AVIATOR’s workflow was guided by both expert knowledge and iterative refinement. We began by interviewing three experienced cybersecurity researchers (including two co-authors) to document how they would manually inject a vulnerability into secure code. Their reasoning consistently followed four stages: (1) understand the program and its dataflow, (2) identify a suitable injection site, such as sanitization checks that prevent the intended flaw, (3) introduce a subtle but plausible modification at this location, for example by weakening or removing a check, and (4) review the resulting change for plausibility and alignment with the intended CWE. These stages provided the blueprint for AVIATOR’s high-level structure.
Each stages was decomposed into elementary subtasks and mapped to specialized agents. For instance, comprehension was assigned to the Code Understanding and Vuln.-Specific Code Understanding agents, and review to a combination of diff-based agents and static analysis. This decomposition reduces the reasoning burden on individual agents and enables the integration of both LLMs and deterministic tools.
The workflow was then refined through repeated experimentation on the SARD test suite 100 (described in §IV-A), which served as our primary development benchmark. Early prototypes revealed systematic shortcomings: the injector often produced truncated code, over-commented edits, or trivial modifications (e.g., whitespace changes). To address these failure modes, we introduced new agents incrementally. Examples include the Polisher, designed to normalize formatting and repair truncated outputs, and the diff-based refinement loop, which detects trivial edits and triggers corrective re-injection.
Crucially, these refinements targeted generic error patterns rather than dataset-specific artifacts. By focusing on issues such as non-substantive edits and truncation, we ensured that improvements generalized across datasets. As shown later in section V, the same workflow behavior transfers to distinct datasets, confirming that the iterative process captured robust error-recovery patterns rather than overfitting to SARD 100.
Through this iterative design, AVIATOR emerged as a workflow organized into two complementary modules (Figure 1). The Vulnerability Injection Module analyzes benign code, identifies potential injection points, and introduces a vulnerability. The Validation Module then checks these candidates using difference analysis, static analysis, and contextual review, providing structured feedback for correction or reinjection. Together, these two modules instantiate the reasoning process elicited from experts: understanding, modifying, reviewing, and revising.
III-C Vulnerability Injection Module
The Vulnerability Injection Module converts a benign function $c^{b}$ into a candidate vulnerable variant $c^{v}$ of a specified CWE. It instantiates the first three stages of the analyst-inspired workflow: comprehension, vulnerability-specific analysis, and transformation. The module comprises the following agents (Figure 1):
III-C1 Code Understanding (LLM)